OpenStudy (anonymous):

Determine if this series converges or diverges: S=ln[1+sin(1/n)/1-sin(1/n)]

OpenStudy (anonymous):

ln[1+sin(1/n)/1-sin(1/n)]=ln[1+(2*sin(1/n)/1-sin(1/n))]~2*sin(1/n)/1-sin(1/n)~2*(1/n)/1=2/n ; then diverges

OpenStudy (anonymous):

Is that correct?

OpenStudy (nenadmatematika):

yes,it's correct...